Circular No. 5661

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

η CARINAE

A. M. van Genderen, Leiden Observatory; M. de Groot, Armagh Observatory; and P. S. The, Astronomical Institute, Amsterdam, write: “A careful inspection of photometry obtained during 1974-1992 in the framework of the Long-Term Photometry of Variables project (Sterken 1983, ESO Messenger No. 33, 10) has revealed the presence of a number of peculiar minima that are at most 0.1 mag deep in the ultraviolet and much less at longer wavelengths. The precise duration of these minima is unknown due to the many time gaps between observations, but it is perhaps on the order of only a few days. A least-squares solution for the 11 observed minima results in an ephemeris with epoch 1989 Apr. 12.59 (± 2.75 m.e.) UT and period 52.32 (± 0.11) days. We tentatively assume that there are no secondary minima, but this is by no means certain. Although this periodicity (if confirmed) suggests an eclipse phenomenon, we urge caution, partly because of the relatively large standard deviation in the O–C values of ± 8.4 days. Bad weather around the predicted minimum on Nov. 10 precluded verification with the 0.50-m Danish telescope at La Silla. Photometry in the uvby system, obtained by D. Beele (Astronomical Institute, University of Münster) before and after this interval, are inconclusive. If the above ephemeris is correct, the dates of the next four minima are 1993 Jan. 1, Feb. 21, Apr. 15, and June 6. The assistance of all observers is requested, especially around these dates. Please contact Arnout van Genderen (Leiden Observatory, Postbus 9513, 2300 RA Leiden, The Netherlands; internet: vangenderen@rulhl1.leidenuniv.nl) as soon as any indication about the validity of the ephemeris has been found. The analysis of the color variations over periods of 1-2 yr shows that these are caused by S Dor-type eruptions similar to those shown by S Dor, AG Car, HR Car and R71 (also observed in the VBLUW system). Therefore, the main object in the core of η Car is almost certainly a hot S Dor-type star or a luminous blue variable (see, e.g., Van Genderen and The 1985, Space Sci. Rev. 39, 317). If the present campaign results in further evidence, the core of η Car could contain an LBV binary.”

NOVA IN THE LARGE MAGELLANIC CLOUD 1992 Further photometry by A. C. Gilmore at Mt. John (cf. IAUC 5659):

Nov. 23.50 UT, V = 12.70 ± 0.03, U–B = –0.75 ± 0.03, B–V = –0.03 ± 0.04, V–R = +0.82 ± 0.04 (reference star CPD −65^{o}433).

1992 November 25 Daniel W. E. Green